# LeanAgent: Lifelong Learning for Formal Theorem Proving

### v2

@@ Abstract
Large Language Models (LLMs) have been successful in mathematical reasoning tasks such as formal theorem proving when integrated with interactive proof assistants like Lean. Existing approaches involve training or fine-tuning an LLM on a specific dataset to perform well on particular domains, such as undergraduate-level mathematics. These methods struggle with generalizability to advanced mathematics. A fundamental limitation is that these approaches operate on static domains, failing to capture how mathematicians often work across multiple domains and projects simultaneously or cyclically. We present LeanAgent, a novel lifelong learning framework for theorem proving that continuously generalizes to and improves on ever-expanding mathematical knowledge without forgetting previously learned knowledge. LeanAgent introduces several key innovations, including a curriculum learning strategy that optimizes the learning trajectory in terms of mathematical difficulty, a dynamic database for efficient management of evolving mathematical knowledge, and progressive training to balance stability and plasticity. LeanAgent successfully proves 162 theorems previously unproved by humans across 23 diverse Lean repositories, many from advanced mathematics. It performs up to 11$\times$ better than the static LLM baseline, proving challenging theorems in domains like abstract algebra and algebraic topology while showcasing a clear progression of learning from basic concepts to advanced topics. In addition, we analyze LeanAgent’s superior performance on key lifelong learning metrics. LeanAgent achieves exceptional scores in stability and backward transfer, where learning new tasks improves performance on previously learned tasks. This emphasizes LeanAgent’s continuous generalizability and improvement, explaining its superior theorem proving performance.
@@ 1 Introduction
Extensive experiments across 23 diverse Lean repositories demonstrate LeanAgent’s advancements in lifelong learning for theorem proving. LeanAgent successfully proves 162 theorems unproven by humans, known as sorry theorems, many from advanced mathematics. For example, it proves difficult sorry theorems from the PFR repository and proves challenging theorems in abstract algebra and algebraic topology related to Coxeter systems and the Hairy Ball Theorem [18, 30]. We find that LeanAgent demonstrates progressive learning in theorem proving, initially proving basic sorry theorems and significantly advancing to more complex ones. It outperforms the static ReProver baseline by up to 11$\times$ in terms of proving new sorry theorems while maintaining performance on known ones.
@@ 3.3 Progressive Training of the Retriever
Although LeanAgent works with any LLM, we provide a specific implementation here. We start with ReProver’s retriever, a fine-tuned version of the ByT5 encoder [79], leveraging its general pre-trained knowledge from mathlib4. We train LeanAgent on the new dataset for an additional epoch. This limited exposure helps prevent overfitting to the new data while allowing LeanAgent to learn essential new information. Before validation, we precompute embeddings for all premises in the corpus to ensure these embeddings are consistent with LeanAgent’s current state. To understand how LeanAgent balances stability and plasticity, we save the model iteration with the highest validation recall for the top ten retrieved premises (R@10). This is a raw plasticity value: it can be used to compute other metrics that describe LeanAgent’s ability to adapt to and understand new types of mathematics in the latest repository (details in Sec. 4). Then, we compute the average test R@10 over all previous datasets the model has progressively trained on, a raw stability value.
@@ 4.1 Experimental Setup
Figure 2: Case studies of LeanAgent’s new proofs. LeanAgent shows a strong understanding of these repositories, often able to retrieve the necessary premises (highlighted) in a way that represents a deep understanding of the mathematics of each repository. For example, LeanAgent proves a sorry theorem from PFR, condRho_of_translate, by simply expanding definitions, showing its deep understanding of the PFR repository. It proved a PFR theorem with just rfl (by reflexivity), demonstrating its understanding of how the terms in the theorem statement interrelate. In addition, LeanAgent could prove re_float from SciLean during lifelong learning, while ReProver could not. Moreover, its MiniF2F proofs demonstrate its ability to generate relatively longer and more complex proofs for complex mathematics. LeanAgent’s proof of wedderburn and invmap.of_eq represents a deep understanding of abstract algebra premises. Impressively, the human proof of wedderburn written after LeanAgent’s proof contains 175 lines, while LeanAgent’s proof is only one line, further demonstrating a deep understanding of abstract algebra.
@@
PFR. LeanAgent can prove a sorry theorem from this cutting-edge repository, while ReProver cannot. It also generalizes to a different commit (not included in progressive training), proving a theorem with just the rfl tactic that ReProver cannot. Interestingly, LeanAgent has a strong enough understanding of logical manipulation within the PFR repository to prove 5 sorry theorems with a $0=1$ placeholder theorem statement (not included in TPPS) (details in Appendix A.5).
@@
Sub-curriculum. In the Formal Book repository, LeanAgent progresses from proving basic real analysis and number theory theorems to mastering advanced abstract algebra, exemplified by its proof of Wedderburn’s Little Theorem. For the Coxeter repository, LeanAgent proves a complex lemma about Coxeter systems, showcasing its proficiency in group theory. In the Hairy Ball Theorem repository, LeanAgent proves a key step of the theorem, demonstrating an understanding of algebraic topology. Only LeanAgent can prove these impressive theorems, demonstrating that it has much more advanced theorem-proving capabilities than ReProver.
@@ 5 Conclusion
We have presented LeanAgent, a lifelong learning framework for theorem proving that achieves continuous generalizability and improvement across diverse mathematical domains. Key components include a curriculum learning strategy, progressive training approach, and custom dynamic database infrastructure. LeanAgent proves 162 sorry theorems across 23 Lean repositories, including from challenging mathematics, highlighting its potential to assist in formalizing complex proofs across multiple domains. For example, LeanAgent successfully proves advanced sorry theorems from the PFR repository and proves challenging theorems in abstract algebra and algebraic topology. It outperforms the ReProver baseline by up to 11$\times$, progressively learning from basic to highly complex mathematical concepts. Moreover, LeanAgent shows significant performance in forgetting measures and backward transfer, achieving a near-perfect composite score of 94%. This explains its continuous generalizability and continuous improvement.
@@ A.5 Further sorry Theorem Proving Discussion
Mathematics in Lean Source. We notice a progression in LeanAgent’s understanding of the Mathematics in Lean Source repository. During lifelong learning, LeanAgent demonstrates a strong grasp of fundamental algebraic structures and basic mathematical operations:
@@
d) Rudimentary Real Analysis: LeanAgent demonstrates an early understanding of properties related to real numbers and absolute values, as shown by C03S05.MyAbs.abs_add, which proves the triangle inequality for real numbers.
10/14 of these proven sorry theorems from Mathematics in Lean Source during the lifelong learning process are from the exercise file for proving identities about algebraic structures. This indicates that LeanAgent starts its understanding of mathematical concepts from the basics.
@@
b) Set Theory and Relations: LeanAgent understands abstract set-theoretic concepts, as shown by C03S01.Subset.trans, which proves that subset relations are transitive.
Now, only 2/7 sorry theorems from Mathematics in Lean Source are from the exercise file for proving identities about algebraic structures. This suggests that lifelong learning allowed LeanAgent to transition to gaining a stronger understanding of how to use premises for more complicated proofs.
We gain some insights from comparing the performance of LeanAgent over time on Mathematics in Lean Source to other setups. For example, the fact that ReProver can handle harder theorems out of the box, such as C03S01.my_lemma3, but fewer theorems overall suggests that it has a broader knowledge base initially but loses performance from a lack of adaptability. Furthermore, Setup 5 proves the same sorry theorems as LeanAgent does during lifelong learning. This suggests that pure curriculum learning without EWC or the Merge All strategy emphasizes understanding easier concepts earlier, showing a learning pattern similar to human progression. This mimics the insights gained from the lifelong learning analysis. However, Setups 3 and 7 (curriculum with EWC and/or Merge All) demonstrate some knowledge plasticity, proving harder theorems during lifelong learning, such as C03S01.Subset.trans in Setup 3. However, this comes at the cost of an understanding of basic theorems, showing catastrophic forgetting. For example, Setups 3 and 7 could not prove the trivial theorems MyGroup.mul_right_inv and MyRing.zero_mul, respectively, during lifelong learning, whereas LeanAgent could. This again aligns with the insights from the composite score from our previous analysis.
By the end of lifelong learning, the sorry theorems that LeanAgent prove from Mathematics in Lean Source are a superset of those that the other setups prove. This shows that LeanAgent’s lifelong learning setup provides continuously improving capabilities to reason about more advanced premises and proofs than other setups. For example, LeanAgent is the only system, except for Setup 6, which can prove theorem C03S05.MyAbs.abs_lt. LeanAgent achieved this through its deep understanding of the available premises, such as abs_lt with a statement similar to C03S05.MyAbs.abs_lt.
@@
It achieves this through its deeper understanding of the le_abs premise, which provides conditions for when an element is less than or equal to the absolute value of another. This suggests that LeanAgent begins by using existing knowledge where possible before trying to deeply understand why existing facts are reasonable.
@@
d) Floating-Point Operations: LeanAgent demonstrates an early understanding of floating-point representations and their correspondence to real numbers, shown by SciLean.re_float, proving that a floating-point number’s real-like part is itself.
@@
b) Sophisticated Bijections: LeanAgent grows in its understanding of product spaces and bijections, proving theorems such as Function.Bijective.Prod.mk.arg_fstsnd.Bijective_rule_simple’, which proves the bijectivity of a function that swaps elements in a product space and Function.Bijective.Equiv.invFun.arg-
@@
, which proves that the composition of a bijection and its inverse remains bijective. Crucially, understanding these theorems might seem simple, but it demonstrates LeanAgent’s understanding of abstract algebraic thinking.
@@
d) Data Structures in Mathematics: LeanAgent understands array types in a mathematical context, proving theorems such as SciLean.ArrayType.ext, which proves that two arrays are equal if their elements are equal at all indices.
It is impressive that LeanAgent could understand a traditionally computer-science-oriented data structure from a mathematical lens, something that some other setups could not understand.
The proofs at this stage are more sophisticated, involving multiple steps and combining various mathematical concepts. This indicates a deeper understanding and ability to connect different areas of mathematics.
@@
Crucially, LeanAgent could prove re_float during lifelong learning while no other setup could. This indicates that LeanAgent’s more measured and stable approach allowed it to understand floating-point representations and their relation to reals. At the same time, other setups prioritized this less with their reduced stability. This may also suggest that continuous improvement allowed LeanAgent to understand new and unique concepts from new repositories.
@@
while LeanAgent could, even during lifelong learning, suggests that this baseline lacks an understanding of foundational concepts. LeanAgent has a better understanding of these concepts from lifelong learning. This is due to the increased stability of LeanAgent and improvement from learning a new task, as shown in the lifelong learning metric analysis in Sec. 4.3.
@@
during lifelong learning. However, as mentioned above, LeanAgent could only prove this theorem at the end of lifelong learning. This suggests that Setup 3 is more plastic during lifelong learning, while LeanAgent remains more stable. This corroborates the analysis from the lifelong learning metrics. However, this again comes at the cost of understanding basic theorems. For example, Setup 3 cannot prove the trivial measure theory theorem SciLean.Measure.prod_volume, whereas LeanAgent could during lifelong learning, again suggesting that it favors plasticity over stability.
@@
LeanAgent suggests that the proof is straightforward after expanding definitions of condRho and considering how rho, the randomness measure, behaves under translation (as captured by the rho_of_translate lemma). The fact that LeanAgent could trivially identify such a short proof using existing premises while the maintainers of the PFR repository did not suggests the power of our approach. This is especially powerful when considering the cutting-edge nature of the content of the PFR repository. This suggests that by understanding the foundations of the PFR lemmas using its improved stability, LeanAgent was able to grasp an understanding of some basic definitions.
@@
Alternate PFR Commit. We also analyze whether LeanAgent can generalize to a different commit of a repository in the curriculum. We choose commit 861715b9bf9482d2442760169cb2a3ff54091f75, because PFR/RhoFunctional.lean, the file from which we proved condRho_of_translate in the newer commit, did not exist in the old commit. This allowed the sorry theorems in the old commit to be more distinct. Impressively, it proved two theorems, including the theorem multiDist_copy about the equality of distributions when copying random variables across different measure spaces that ReProver could not. LeanAgent achieved this with just the tactic rfl, indicating the theorem statements are true by reflexivity.
This suggests that LeanAgent has a deep enough understanding of PFR, even beyond the commit it learned from, to notice the connection to reflexivity that even humans did not notice at this commit.
This commit also provides an interesting example of LeanAgent’s power. The PFR maintainers used 0 = 1 as a placeholder for some sorry theorems, five of which are multiTau_min_exists, multiTau_min_sum_le, sub_multiDistance_le, sub_condMultiDistance_le, sub_condMultiDistance_le’. Interestingly, LeanAgent is powerful enough to find loopholes and prove these theorems with this proof:
It combines the known fact that $0\neq 1$ with the placeholder theorem multidist_eq_zero, which states $0=1$. As such, it proves False, allowing it to derive anything, including $0=1$. This demonstrates LeanAgent’s deep understanding of logical manipulation within the PFR repository.
@@
a) Basic Arithmetic and Number Theory: ReProver could handle simple arithmetic and modular arithmetic problems, such as mathd_numbertheory_254, a theorem about modular arithmetic and basic addition, mathd_numbertheory_342, a theorem about basic divisibility, and mathd_algebra_304, a theorem about simple exponentiation. These proofs only rely on the norm_num tactic, which evaluates arithmetic expressions. This suggests a less sophisticated understanding of mathematics at the start of lifelong learning.
@@
a) Advanced Number Theory: LeanAgent showed a more advanced understanding of number theory, proving theorems like mathd_numbertheory_293, a complex theorem about divisibility involving a complex expression and mathd_numbertheory_233, a theorem dealing with modular arithmetic in $\text{ZMod}(11^{2})$.
@@
e) Complex Quantifiers and Inequalities: LeanAgent increased its understanding of more complex logical statements, such as amc12a_2002_p6, a theorem involving multiple existential quantifiers and inequalities.
The proofs at this later stage are more sophisticated, usually involving multiple steps and combining various mathematical concepts or indicating a deeper understanding and ability to connect different areas of mathematics, mirroring the progression observed in Mathematics in Lean Source and SciLean. For example, as shown above, LeanAgent provides a one-line proof to the relatively advanced theorem mathd_numbertheory_233. The proof means the hypothesis directly proves the goal. This suggests that LeanAgent has developed a deep understanding of modular arithmetic and can recognize when a given hypothesis is sufficient to prove the goal without additional steps.
Furthermore, as shown above, LeanAgent uses four tactics to prove the induction_12dvd4expnp1p20 theorem. This demonstrates its ability to handle more complex number theory proofs and use advanced tactics. This again shows that LeanAgent can recognize when its understanding its deep enough to not require complex premise retrieval.
LeanAgent demonstrates a similar understanding of the theorem amc12a_2002_p6. Notably, it combines the simple premises lt_add_of_pos_right, which describes how an element is less than that element added with a positive one, and zero_lt_one, which states that 0 is less than 1, with more advanced tactics like lift, cases, and exact with complex term construction. This demonstrates its ability to reuse foundational understanding for more complex proofs of abstract mathematical concepts, showing its stability.
@@
LeanAgent’s proof of the wedderburn theorem impressively represents a deep understanding of algebraic structures. By using the Field.toIsField premise, LeanAgent shows that it has grasped what makes a ring a field and can apply this knowledge efficiently. This requires a comprehensive understanding of ring theory and field properties. Impressively, the human proof of wedderburn written after LeanAgent’s proof contains 175 lines, while LeanAgent’s proof is only one line. This suggests that LeanAgent has developed a deep, intuitive grasp of abstract algebra.
Coxeter. LeanAgent could not prove sorry theorems from the Coxeter repository during lifelong learning. However, by the end of lifelong learning, LeanAgent demonstrates growing proficiency in more complex algebraic structures, proving the lemma invmap.of_eq about Coxeter systems, again showing the ability to work with advanced concepts in group theory and abstract algebra. This corroborates the deeper understanding of abstract algebra necessary to prove the wedderburn theorem.
@@
Hairy Ball Theorem. Moreover, LeanAgent could not prove sorry theorems from the Hairy Ball Theorem repository during lifelong learning. However, at the end of lifelong learning, LeanAgent again demonstrates growing proficiency in algebraic topology. It proves HairyBallDiff, which states a key step in the Hairy Ball Theorem, demonstrating an understanding of vector spaces, norms, and their connections to topological concepts.

### v4

@@ Abstract
Large Language Models (LLMs) have been successful in mathematical reasoning tasks such as formal theorem proving when integrated with interactive proof assistants like Lean. Existing approaches involve training or fine-tuning an LLM on a specific dataset to perform well on particular domains, such as undergraduate-level mathematics. These methods struggle with generalizability to advanced mathematics. A fundamental limitation is that these approaches operate on static domains, failing to capture how mathematicians often work across multiple domains and projects simultaneously or cyclically. We present LeanAgent, a novel lifelong learning framework for theorem proving that continuously generalizes to and improves on ever-expanding mathematical knowledge without forgetting previously learned knowledge. LeanAgent introduces several key innovations, including a curriculum learning strategy that optimizes the learning trajectory in terms of mathematical difficulty, a dynamic database for efficient management of evolving mathematical knowledge, and progressive training to balance stability and plasticity. LeanAgent successfully proves 162 theorems previously unproved by humans across 23 diverse Lean repositories, many from advanced mathematics. It performs significantly better than the static LLM baseline, proving challenging theorems in domains like abstract algebra and algebraic topology while showcasing a clear progression of learning from basic concepts to advanced topics. In addition, we analyze LeanAgent’s superior performance on key lifelong learning metrics. LeanAgent achieves exceptional scores in stability and backward transfer, where learning new tasks improves performance on previously learned tasks. This emphasizes LeanAgent’s continuous generalizability and improvement, explaining its superior theorem-proving performance.
@@ 1 Introduction
Extensive experiments across 23 diverse Lean repositories demonstrate LeanAgent’s advancements in lifelong learning for theorem proving. LeanAgent successfully proves 162 theorems unproven by humans, known as sorry theorems, many from advanced mathematics. This count includes mathematical proofs and exploits found within Lean’s type system. For example, it proves difficult sorry theorems from the PFR repository and proves challenging theorems in abstract algebra and algebraic topology related to Coxeter systems and the Hairy Ball Theorem [18, 30]. We find that LeanAgent demonstrates progressive learning in theorem proving, initially proving basic sorry theorems and significantly advancing to more complex ones. It significantly outperforms the static ReProver baseline in terms of proving new sorry theorems while maintaining performance on known ones. We have issued pull requests to the respective repositories with the newly proven sorry theorems. Some of them turned out to be exploits due to the loopholes in the repositories that are being fixed, and in future, should no longer be an issue in the future after the fixes. We plan to run LeanAgent periodically on all the repositories as they are being updated with human-written proofs, enabling LeanAgent to continue learning and prove more theorems over time.
@@ 3.3 Progressive Training of the Retriever
Although LeanAgent works with any LLM, we provide a specific implementation here. We start with ReProver’s retriever, a fine-tuned version of the ByT5 encoder [79], leveraging its general pre-trained knowledge from mathlib4. We train LeanAgent on the new dataset for an additional epoch. This limited exposure helps prevent overfitting to the new data while allowing LeanAgent to learn essential new information. Before validation, we precompute embeddings for all premises in the corpus to ensure these embeddings are consistent with LeanAgent’s current state. To understand how LeanAgent balances stability and plasticity, we save the model iteration with the highest validation recall for the top ten retrieved premises (R@10). This is a raw plasticity value: it can be used to compute other metrics that describe LeanAgent’s ability to adapt to and handle new types of mathematics in the latest repository (details in Sec. 4). Then, we compute the average test R@10 over all previous datasets the model has progressively trained on, a raw stability value.
@@ 4.1 Experimental Setup
Figure 2: Case studies of LeanAgent’s new proofs. LeanAgent shows an ability to work with these repositories, often able to retrieve the necessary premises (highlighted). For example, LeanAgent proves a sorry theorem from PFR, condRho_of_translate, by simply expanding definitions, showing its proving ability on the PFR repository. It proved a PFR theorem with just rfl (by reflexivity), demonstrating its grasp of how the terms in the theorem statement interrelate. In addition, LeanAgent could prove re_float from SciLean during lifelong learning, while ReProver could not. Moreover, its MiniF2F proofs demonstrate its ability to generate relatively longer and more complex proofs for complex mathematics. LeanAgent’s proof of wedderburn and invmap.of_eq represents its theorem proving capabilities with abstract algebra premises. Impressively, the human proof of wedderburn written after LeanAgent’s proof contains 175 lines, while LeanAgent’s proof is only one line, further demonstrating its performance in abstract algebra.
@@
PFR. LeanAgent can prove a sorry theorem from this cutting-edge repository, while ReProver cannot. It also generalizes to a different commit (not included in progressive training), proving a theorem with just the rfl tactic that ReProver cannot. Interestingly, LeanAgent has a strong enough ability of logical manipulation within the PFR repository to prove 5 sorry theorems with a $0=1$ placeholder theorem statement (not included in TPPS) (details in Appendix A.5).
@@
Sub-curriculum. In the Formal Book repository, LeanAgent progresses from proving basic real analysis and number theory theorems to mastering advanced abstract algebra, exemplified by its proof of Wedderburn’s Little Theorem. For the Coxeter repository, LeanAgent proves a complex lemma about Coxeter systems, showcasing its proficiency in group theory. In the Hairy Ball Theorem repository, LeanAgent proves a key step of the theorem, demonstrating performance in algebraic topology. Only LeanAgent can prove these impressive theorems, demonstrating that it has much more advanced theorem-proving capabilities than ReProver.
Limitations of TPPS: There is no standardized measure of proof difficulty and proving performance, especially in a life-long learning setting. TPPS was an attempt towards this. The core idea is to have a higher reward for newly proven theorems, which aligns with the objective of lifelong learning. In future works, we also plan to investigate some more sophisticated measures that reward not only newly proved theorems but also difficult ones. However, TPPS has several important limitations. It is vulnerable to artificial inflation through exploitation of technicalities, does not account for varying theorem difficulties, and lacks standardization across mathematical domains. The metric focuses on quantity and faces challenges in ensuring comparisons across diverse repositories. These shortcomings highlight the need for more sophisticated evaluation frameworks in theorem proving, especially for lifelong learning contexts. Future work should focus on developing standardized, robust metrics that address these issues and provide more meaningful assessments of theorem-proving capabilities.
@@ 5 Conclusion
We have presented LeanAgent, a lifelong learning framework for theorem proving that achieves continuous generalizability and improvement across diverse mathematical domains. Key components include a curriculum learning strategy, progressive training approach, and custom dynamic database infrastructure. LeanAgent proves 162 sorry theorems across 23 Lean repositories, including from challenging mathematics, highlighting its potential to assist in formalizing complex proofs across multiple domains and identifying system exploits. For example, LeanAgent successfully proves advanced sorry theorems from the PFR repository and proves challenging theorems in abstract algebra and algebraic topology. It outperforms the ReProver baseline by up to 11$\times$, progressively learning from basic to highly complex mathematical concepts. Moreover, LeanAgent shows significant performance in forgetting measures and backward transfer, achieving a near-perfect composite score of 94%. This explains its continuous generalizability and continuous improvement.
@@ A.2 Experiment Implementation Details
Moreover, we do not compare with Aesop because it is not an ML model. We aim to improve upon ML research for theorem proving, such as ReProver. Moreover, Aesop is not a framework, but ReProver was included as the starting point of the LeanAgent framework, which is why we compare LeanAgent to ReProver. Rather than comparing against existing tools, we aim to understand how lifelong learning can work in theorem proving.
@@ A.5 Further sorry Theorem Proving Discussion
Mathematics in Lean Source. We notice a progression in LeanAgent’s proving ability in the Mathematics in Lean Source repository. During lifelong learning, LeanAgent demonstrates a strong grasp of fundamental algebraic structures and basic mathematical operations:
@@
d) Rudimentary Real Analysis: LeanAgent demonstrates an early capability to handle properties related to real numbers and absolute values, as shown by C03S05.MyAbs.abs_add, which proves the triangle inequality for real numbers.
10/14 of these proven sorry theorems from Mathematics in Lean Source during the lifelong learning process are from the exercise file for proving identities about algebraic structures. This indicates that LeanAgent starts its grasp of mathematical concepts from the basics.
@@
b) Set Theory and Relations: LeanAgent handles abstract set-theoretic concepts, as shown by C03S01.Subset.trans, which proves that subset relations are transitive.
Now, only 2/7 sorry theorems from Mathematics in Lean Source are from the exercise file for proving identities about algebraic structures. This suggests that lifelong learning allowed LeanAgent to transition to gaining a stronger ability to work with premises for more complicated proofs.
We gain some insights from comparing the performance of LeanAgent over time on Mathematics in Lean Source to other setups. For example, the fact that ReProver can handle harder theorems out of the box, such as C03S01.my_lemma3, but fewer theorems overall suggests that it has a broader knowledge base initially but loses performance from a lack of adaptability. Furthermore, Setup 5 proves the same sorry theorems as LeanAgent does during lifelong learning. This suggests that pure curriculum learning without EWC or the Merge All strategy emphasizes grasping easier concepts earlier, showing a learning pattern similar to human progression. This mimics the insights gained from the lifelong learning analysis. However, Setups 3 and 7 (curriculum with EWC and/or Merge All) demonstrate some knowledge plasticity, proving harder theorems during lifelong learning, such as C03S01.Subset.trans in Setup 3. However, this comes at the cost of proving basic theorems, showing catastrophic forgetting. For example, Setups 3 and 7 could not prove the trivial theorems MyGroup.mul_right_inv and MyRing.zero_mul, respectively, during lifelong learning, whereas LeanAgent could. This again aligns with the insights from the composite score from our previous analysis.
By the end of lifelong learning, the sorry theorems that LeanAgent prove from Mathematics in Lean Source are a superset of those that the other setups prove. This shows that LeanAgent’s lifelong learning setup provides continuously improving capabilities to reason about more advanced premises and proofs than other setups. For example, LeanAgent is the only system, except for Setup 6, which can prove theorem C03S05.MyAbs.abs_lt. LeanAgent achieved this by using the available premises, such as abs_lt with a statement similar to C03S05.MyAbs.abs_lt.
@@
It achieves this through its use of the le_abs premise, which provides conditions for when an element is less than or equal to the absolute value of another. This suggests that LeanAgent begins by using existing knowledge where possible before trying to realizing why existing facts are reasonable.
@@
d) Floating-Point Operations: LeanAgent demonstrates an early grasp of floating-point representations and their correspondence to real numbers, shown by SciLean.re_float, proving that a floating-point number’s real-like part is itself.
@@
b) Sophisticated Bijections: LeanAgent grows in its ability to work with product spaces and bijections, proving theorems such as Function.Bijective.Prod.mk.arg_fstsnd.Bijective_rule_simple’, which proves the bijectivity of a function that swaps elements in a product space and Function.Bijective.Equiv.invFun.arg-
@@
, which proves that the composition of a bijection and its inverse remains bijective. Crucially, these theorems might seem simple, but they demonstrate LeanAgent’s capability to handle of abstract algebraic thinking.
@@
d) Data Structures in Mathematics: LeanAgent navigates proofs involving array types in a mathematical context, proving theorems such as SciLean.ArrayType.ext, which proves that two arrays are equal if their elements are equal at all indices.
It is impressive that LeanAgent could prove a theorem about a traditionally computer-science-oriented data structure from a mathematical lens, which some other setups could not do.
The proofs at this stage are more sophisticated, involving multiple steps and combining various mathematical concepts. This indicates a deeper ability to connect different areas of mathematics.
@@
Crucially, LeanAgent could prove re_float during lifelong learning while no other setup could. This indicates that LeanAgent’s more measured and stable approach allowed it to grasp floating-point representations and their relation to reals. At the same time, other setups prioritized this less with their reduced stability. This may also suggest that continuous improvement allowed LeanAgent to process new and unique concepts from new repositories.
@@
while LeanAgent could, even during lifelong learning, suggests that this baseline cannot handle foundational concepts. LeanAgent has a better grasp of these concepts from lifelong learning. This is due to the increased stability of LeanAgent and improvement from learning a new task, as shown in the lifelong learning metric analysis in Sec. 4.3.
@@
during lifelong learning. However, as mentioned above, LeanAgent could only prove this theorem at the end of lifelong learning. This suggests that Setup 3 is more plastic during lifelong learning, while LeanAgent remains more stable. This corroborates the analysis from the lifelong learning metrics. However, this again comes at the cost of grasping basic theorems. For example, Setup 3 cannot prove the trivial measure theory theorem SciLean.Measure.prod_volume, whereas LeanAgent could during lifelong learning, again suggesting that it favors plasticity over stability.
@@
LeanAgent suggests that the proof is straightforward after expanding definitions of condRho and considering how rho, the randomness measure, behaves under translation (as captured by the rho_of_translate lemma). The fact that LeanAgent could trivially identify such a short proof using existing premises while the maintainers of the PFR repository did not suggests the power of our approach. This is especially powerful when considering the cutting-edge nature of the content of the PFR repository. This suggests that by learning the foundations of the PFR lemmas using its improved stability, LeanAgent was able to grasp some basic definitions.
@@
Alternate PFR Commit. We also analyze whether LeanAgent can generalize to a different commit of a repository in the curriculum. We choose commit 861715b9bf9482d2442760169cb2a3ff54091f75, because PFR/RhoFunctional.lean, the file from which we proved condRho_of_translate in the newer commit, did not exist in the old commit. This allowed the sorry theorems in the old commit to be more distinct. It proved two theorems, including the theorem multiDist_copy about the equality of distributions when copying random variables across different measure spaces that ReProver could not. LeanAgent achieved this with just the tactic rfl. However, these statements were about multiDist, another sorry theorem. Since any two instances of sorry are automatically equal by rfl, LeanAgent exploits this technicality to prove the two theorems.
This commit also provides another interesting example. The PFR maintainers used 0 = 1 as a placeholder for some sorry theorems, five of which are multiTau_min_exists, multiTau_min_sum_le, sub_multiDistance_le, sub_condMultiDistance_le, sub_condMultiDistance_le’. Interestingly, LeanAgent finds loopholes and proves these theorems with this proof:
It combines the known fact that $0\neq 1$ with the placeholder theorem multidist_eq_zero, which states $0=1$. As such, it proves False, allowing it to derive anything, including $0=1$.
@@
a) Basic Arithmetic and Number Theory: ReProver could handle simple arithmetic and modular arithmetic problems, such as mathd_numbertheory_254, a theorem about modular arithmetic and basic addition, mathd_numbertheory_342, a theorem about basic divisibility, and mathd_algebra_304, a theorem about simple exponentiation. These proofs only rely on the norm_num tactic, which evaluates arithmetic expressions. This suggests a less sophisticated proving ability of mathematics at the start of lifelong learning.
@@
a) Advanced Number Theory: LeanAgent showed a more advanced grasp of number theory, proving theorems like mathd_numbertheory_293, a complex theorem about divisibility involving a complex expression and mathd_numbertheory_233, a theorem dealing with modular arithmetic in $\text{ZMod}(11^{2})$.
@@
e) Complex Quantifiers and Inequalities: LeanAgent increased its ability to prove more complex logical statements, such as amc12a_2002_p6, a theorem involving multiple existential quantifiers and inequalities.
The proofs at this later stage are more sophisticated, usually involving multiple steps and combining various mathematical concepts or indicating a better ability to connect different areas of mathematics, mirroring the progression observed in Mathematics in Lean Source and SciLean. For example, as shown above, LeanAgent provides a one-line proof to the relatively advanced theorem mathd_numbertheory_233. The proof means the hypothesis directly proves the goal. This suggests that LeanAgent grasps modular arithmetic and recognize when a given hypothesis is sufficient to prove the goal without additional steps.
Furthermore, as shown above, LeanAgent uses four tactics to prove the induction_12dvd4expnp1p20 theorem. This demonstrates its ability to handle more complex number theory proofs and use advanced tactics. This again shows that LeanAgent can recognize when it does not require complex premise retrieval.
LeanAgent demonstrates a similar grasp of the theorem amc12a_2002_p6. Notably, it combines the simple premises lt_add_of_pos_right, which describes how an element is less than that element added with a positive one, and zero_lt_one, which states that 0 is less than 1, with more advanced tactics like lift, cases, and exact with complex term construction. This demonstrates its ability to reuse foundational concepts for more complex proofs of abstract mathematical concepts, showing its stability.
@@
LeanAgent’s proof of the wedderburn theorem represents the ability to handle algebraic structures. By using the Field.toIsField premise, LeanAgent shows that it has grasped what makes a ring a field and can apply this knowledge efficiently. This requires a comprehensive grasp of ring theory and field properties. Impressively, the human proof of wedderburn written after LeanAgent’s proof contains 175 lines, while LeanAgent’s proof is only one line. This suggests that LeanAgent has developed a grasp of abstract algebra.
Coxeter. LeanAgent could not prove sorry theorems from the Coxeter repository during lifelong learning. However, by the end of lifelong learning, LeanAgent demonstrates growing proficiency in more complex algebraic structures, proving the lemma invmap.of_eq about Coxeter systems, again showing the ability to work with advanced concepts in group theory and abstract algebra. This corroborates the handling of abstract algebra necessary to prove the wedderburn theorem.
@@
Hairy Ball Theorem. Moreover, LeanAgent could not prove sorry theorems from the Hairy Ball Theorem repository during lifelong learning. However, at the end of lifelong learning, LeanAgent again demonstrates growing proficiency in algebraic topology. It proves HairyBallDiff, which states a key step in the Hairy Ball Theorem, demonstrating a grasp of vector spaces, norms, and their connections to topological concepts.